Can you show the work on the following equation so I can "see how" it is done?
x^2 -6x -16 = 0
x^2 means x squared.
Thank you!!

Answer by jorel1380(3719)

x2-6x-16=0
(x+2)(x-8)=0
x=-2 or 8..